Suppose the market demand and supply equations for as product are
Qd = 25-3p
Qs=10+2p
What are the equilibrium price quantity for this product?
1
Expert's answer
2020-09-07T11:12:18-0400

At equilibrium, Quantity demand is equal to the quantity supplied. Therefore,

Qd=Qs

Solving mathematically

25-3p=10+2p

25-10=2p+3p

15=5p thus p= 3

Hence we can substitute in Quantity demand function as follows

25-3p=25-3(3)

=16

Equilibrium price quantity is 16